An asteroid has an momentum of 5.6E6 kg·m/s in the +x direction and a momentum of...

  1. An asteroid has an momentum of 5.6E6 kg·m/s in the +x direction and a momentum of 3.7E6 kg·m/s in the +y direction. What is the angle, in degrees, of the resulting momentum?   

  2. A 4.0 kg ball with a velocity of 1.4 m/s and a 2.5 kg ball with a velocity of -1.0 m/s collide. If the collision was perfectly elastic and the velocity of the 4.0 kg ball was -0.44615 m/s (extra digits kept to avoid round-off error in the final answer) after the collision, what was the velocity of the 2.5 kg ball? Hint: Since the velocity of one of the balls after the collision is provided, Conservation of Momentum (p = po) may be applied directly. Assume one-dimensional motion, i.e., all vectors and movements occur along a line. Assume no external forces.   
Answer #1

3.

x = arctan (3.7 /5.6)

x = 33.435

=====

4.

using momentum conservation

m1 u1 + m2 u2 = m1 v1 + m2 v2

4* 1.4 - 2.5* 1 = 4*(-0.44615) + 2.5 v2

v2 = 1.90584 m/s

=====

Comment before rate in case any doubt, will reply for sure.. goodluck